# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ 3.3.2 Using scaled cluster-period totals
Define $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{T}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,\allowbreak r_{iJ\cdot}(a))^{T}$. Under Assumptions 1 and 2, and regularity conditions (C1), (C6) and (C7) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $I^{-1}\sum_{i}\widetilde{Y}_{ij\cdot}(a)^{4}=O(1)$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
Comparing results in Theorems 5 and 6, for marginal variances, we find that covariate adjustment always improves the asymptotic efficiency because $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}(a,a^{\prime})\}$ from the regressions using scaled cluster-period totals. Theorems 5 and 6 also provide the theoretical foundation for using the HC variance estimator for variance estimation in SR-IREs, which was pointed out in Roth and Sant’Anna (2023) without a theoretical proof. Finally, the ANCOVA estimator $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})$ obtained via the OLS fit of $\widetilde{Y}_{ij\cdot}$ on $(\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ij}^{T},\overrightarrow{\mathbbm{1}}_{ij}^{T}\otimes\bm{C}_{ij}^{c})$ is less efficient than $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})$ and may not improve the asymptotic efficiency of $\widehat{\tau}_{j,\mathrm{T}}(a,a^{\prime})$, as in simpler cases under parallel-arm CREs (Su and Ding, 2021). Similar $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})$ and $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$, $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})$ is also consistent and asymptotically normal for $\bm{\tau}(a,a^{\prime})$, and the associated HC variance estimator $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is conservative for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})\}$, in the Löwner ordering.
@@ 3.4 Generalization to summary estimands
The summary estimand in (5) is a weighted sum of $\tau_{j}(a,a^{\prime})$ for $j=1,\ldots,J$, assuming $a<a^{\prime}\in\mathcal{A}$. We can estimate $\theta$ via plugging in estimates of $\tau_{j}(a,a^{\prime})$ from any proposed regression estimators, e.g., $\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}=\sum_{j=1}^{J}\sum_{a<a^{\prime}}b_{\theta,j}(a,a^{\prime})\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ if we select the regression estimator using individual-level data with covariate adjustment. The variance of $\widehat{\theta}$, $\mathrm{se}^{2}(\widehat{\theta})=\bm{b}_{\theta}^{T}\mathbf{Cov}(\widehat{\bm{\tau}})\bm{b}_{\theta}$, where $\bm{b}_{\theta}=(\bm{b}_{\theta}(1,2),\ldots,\bm{b}_{\theta}(1,\infty)^{T},\bm{b}_{\theta}(2,3)^{T},\ldots,\allowbreak\bm{b}_{\theta}(2,\infty)^{T},\ldots,\allowbreak\bm{b}_{\theta}(J,\infty)^{T})^{T}$ with $\bm{b}_{\theta}(a,a^{\prime})=(b_{\theta,1}(a,a^{\prime}),\allowbreak\ldots,b_{\theta,j}(a,a^{\prime}))^{T}$ for $a,a\in\mathcal{A}$ and $a<a^{\prime}$, can be estimated using the CR or HC variance estimator depending on the selected regression estimator. Continuing with the example of $\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}$, we have the CR variance estimator as $\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}})=\bm{b}_{\theta}^{T}\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})\bm{b}_{\theta}$, where $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{T}$ with
$\bm{Q}=\begin{pmatrix}\bm{Q}(1,2)^{T}&\cdots&\bm{Q}(1,\infty)^{T}&\bm{Q}(2,3)^{T}&\cdots&\bm{Q}(2,\infty)^{T}&\cdots&\bm{Q}(J,\infty)^{T}\end{pmatrix}^{T}.$
Table 2 summarizes regression estimators and their corresponding covariance estimators.
@@
$\widehat{\bm{\tau}}_{\mathrm{I}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}})=\bm{Q}\bm{E}_{\mathrm{I}}\bm{Q}^{T}$ $\widehat{\bm{\tau}}_{\mathrm{A}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{A}})=\bm{Q}\bm{E}_{\mathrm{A}}\bm{Q}^{T}$
$\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{T}$ $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{A}})=\bm{Q}\bm{E}_{\mathrm{A},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{T}$
$\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{ancova}}\bm{Q}^{T}$ $\widehat{\bm{\tau}}_{\mathrm{T}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{T}})=\bm{Q}\bm{E}_{\mathrm{T}}\bm{Q}^{T}$
$\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{T}})=\bm{Q}\bm{E}_{\mathrm{T},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{T}$
[…]
$\bm{Q}=\begin{pmatrix}\bm{Q}_{[1]}&\cdots&\bm{Q}_{[J]}&\bm{Q}_{[\infty]}\end{pmatrix}$,
where $\bm{Q}_{[a]}$ for $a\in\mathcal{A}$ is a matrix of dimensions $J^{2}(J+1)/2\times J$, e.g.,
[…]
The quantities above enable us to define
$\bm{V}_{c}(\bm{\xi}_{i\cdot})=(I-1)\sum_{a\in\mathcal{A}}\frac{1}{I(a)}\bm{Q}_{[a]}\bm{S}^{2}(\bm{\xi}_{i\cdot},a)\bm{Q}_{[a]}^{T}$
and
$\bm{V}(\bm{\xi}_{i\cdot})=\bm{V}_{c}(\bm{\xi}_{i\cdot})-(I-1)I^{-1}\left\{\sum_{a\in\mathcal{A}}\bm{Q}_{[a]}\bm{S}^{2}(\bm{\xi}_{i\cdot},a)\bm{Q}_{[a]}^{T}+\sum_{a\neq a^{\prime}}\bm{Q}_{[a]}\bm{S}^{2}(\bm{\xi}_{i\cdot},a,a^{\prime})\bm{Q}_{[a^{\prime}]}^{T}\right\},$
[…]
Theorem 7
  1. (a)
    Under conditions given in Theorem 1 with $\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{I}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{I}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}(\widehat{\theta}_{\mathrm{I}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\widetilde{\bm{\epsilon}}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  2. (b)
    Let $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under conditions given in Theorem 2 with $\bm{V}(\bm{r}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\bm{r}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\bm{r}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  3. (c)
    Let $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\sum_{a\in\mathcal{A}}q(a)\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under conditions given in Corollary 1 with $\bm{V}(\bm{r}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, for $a,a^{\prime}\in\mathcal{A}$ and $a\neq a^{\prime}$, $\widehat{\theta}_{\mathrm{I}}^{\mathrm{ancova}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{ancova}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\bm{r}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{ancova}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\bm{r}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  4. (d)
    Under conditions given in Theorem 3 with $\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{A}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{A}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{HC}}^{2}(\widehat{\theta}_{\mathrm{A}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\widetilde{\bm{\epsilon}}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  5. (e)
    Let $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{A}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under conditions given in Theorem 4 with $\bm{V}(\bm{r}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\bm{r}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{HC}}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\bm{r}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  6. (f)
    Let $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under conditions given in Theorem 5 with $\bm{V}(\bm{r}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{T}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{T}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\bm{r}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{HC}}^{2}(\widehat{\theta}_{\mathrm{T}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\bm{r}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
  7. (g)
    Let $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{T}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under conditions given in Theorem 6 with $\bm{V}(\bm{r}_{i\cdot})\nrightarrow\bm{0}_{J^{2}(J+1)/2\times J^{2}(J+1)/2}$, $\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}=\theta+o_{\mathbb{P}}(1)$, $I^{1/2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}-\theta)\xrightarrow{d}\mathcal{N}(0,\bm{b}_{\theta}^{T}\bm{V}(\bm{r}_{i\cdot})\bm{b}_{\theta})$, and $I\times\widehat{\mathrm{se}}_{\mathrm{HC}}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}})=\bm{b}_{\theta}^{T}\bm{V}_{c}(\bm{r}_{i\cdot})\bm{b}_{\theta}+o_{\mathbb{P}}(1)$.
[…]
$\bm{V}(\bm{\xi}_{i\cdot})=\bm{V}_{c}(\bm{\xi}_{i\cdot})-I^{-1}\left\{\sum_{i=1}^{I}\sum_{a\in\mathcal{A}}\bm{Q}_{[a]}\bm{\xi}_{i\cdot}(a)\right\}\left\{\sum_{i=1}^{I}\sum_{a\in\mathcal{A}}\bm{Q}_{[a]}\bm{\xi}_{i\cdot}(a)\right\}^{T},$
where the second term on the right-hand side is generally non-negative definite and inestimable, and the CR and HC variance estimators are consistent for $\bm{V}_{c}(\bm{\xi}_{i\cdot})$.

### v3

@@ 3.3.2 Using scaled cluster-period totals
Define $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{T}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,\allowbreak r_{iJ\cdot}(a))^{\top}$. Under Assumptions 1 and 2, and regularity conditions (C1), (C6), and (C7) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $I^{-1}\sum_{i}\widetilde{Y}_{ij\cdot}(a)^{4}=O(1)$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
Comparing results in Theorems 5 and 6, for marginal variances, we find that covariate adjustment always improves the asymptotic efficiency because $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}(a,a^{\prime})\}$ from the regressions using scaled cluster-period totals. Theorems 5 and 6 also provide the theoretical foundation for using the HC variance estimator for variance estimation in SR-IREs, which was pointed out in Roth and Sant’Anna (2023) without a theoretical proof. Finally, the ANCOVA estimator $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})$ obtained via the OLS fit of $\widetilde{Y}_{ij\cdot}$ on $(\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ij}^{\top},\boldsymbol{\mathbbm{1}}_{ij}^{\top}\otimes\bm{C}_{ij}^{c})$ is less efficient than $\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})$ and may not improve the asymptotic efficiency of $\widehat{\tau}_{j,\mathrm{T}}(a,a^{\prime})$, as in simpler cases under parallel-arm CREs (Su and Ding, 2021). Similar $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})$ and $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$, $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})$ is also consistent and asymptotically normal for $\bm{\tau}(a,a^{\prime})$, and the associated HC variance estimator $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is conservative for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{ancova}}(a,a^{\prime})\}$, in the Löwner ordering.
@@ 3.4 Generalization to summary estimands
The summary estimand in (5) is a weighted sum of $\tau_{j}(a,a^{\prime})$ for $j=1,\ldots,J$, assuming $a<a^{\prime}\in\mathcal{A}$. We can estimate $\theta$ via plugging in estimates of $\tau_{j}(a,a^{\prime})$ from any proposed regression estimators, e.g., $\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}=\sum_{j=1}^{J}\sum_{a<a^{\prime}}b_{\theta,j}(a,a^{\prime})\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ if we select the regression estimator using individual-level data with covariate adjustment. The variance of $\widehat{\theta}$, $\mathrm{se}^{2}(\widehat{\theta})=\bm{b}_{\theta}^{\top}\mathbf{Cov}(\widehat{\bm{\tau}})\bm{b}_{\theta}$, where $\bm{b}_{\theta}=(\bm{b}_{\theta}(1,2),\ldots,\bm{b}_{\theta}(1,\infty)^{\top},\bm{b}_{\theta}(2,3)^{\top},\ldots,\allowbreak\bm{b}_{\theta}(2,\infty)^{\top},\ldots,\allowbreak\bm{b}_{\theta}(J,\infty)^{\top})^{\top}$ with $\bm{b}_{\theta}(a,a^{\prime})=(b_{\theta,1}(a,a^{\prime}),\allowbreak\ldots,b_{\theta,j}(a,a^{\prime}))^{\top}$ for $a,a\in\mathcal{A}$ and $a<a^{\prime}$, can be estimated using the CR or HC variance estimator depending on the selected regression estimator. Continuing with the example of $\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}}$, we have the CR variance estimator as $\widehat{\mathrm{se}}_{\mathrm{CR}}^{2}(\widehat{\theta}_{\mathrm{I}}^{\mathrm{adj}})=\bm{b}_{\theta}^{\top}\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})\bm{b}_{\theta}$, where $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{\top}$ with
$\bm{Q}=\begin{pmatrix}\bm{Q}(1,2)^{\top}&\cdots&\bm{Q}(1,\infty)^{\top}&\bm{Q}(2,3)^{\top}&\cdots&\bm{Q}(2,\infty)^{\top}&\cdots&\bm{Q}(J,\infty)^{\top}\end{pmatrix}^{\top}$.
[…]
@@
$\widehat{\bm{\tau}}_{\mathrm{I}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}})=\bm{Q}\bm{E}_{\mathrm{I}}\bm{Q}^{\top}$ $\widehat{\bm{\tau}}_{\mathrm{A}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{A}})=\bm{Q}\bm{E}_{\mathrm{A}}\bm{Q}^{\top}$
$\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{\top}$ $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{A}})=\bm{Q}\bm{E}_{\mathrm{A},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{\top}$
$\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}$ $\widehat{\mathbf{Cov}}_{\mathrm{CR}}(\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}})=\bm{Q}\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{ancova}}\bm{Q}^{\top}$ $\widehat{\bm{\tau}}_{\mathrm{T}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{T}})=\bm{Q}\bm{E}_{\mathrm{T}}\bm{Q}^{\top}$
$\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}$ $\widehat{\mathbf{Cov}}_{\mathrm{HC}}(\widehat{\bm{\tau}}_{\mathrm{T}})=\bm{Q}\bm{E}_{\mathrm{T},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}^{\top}$